OpenStudy (anonymous):

The least possible fraction that is greater than 1/2 ?? The least possible fraction that is greater than 1/2 ?? @Mathematics

OpenStudy (jamesj):

You want the smallest fraction p/q such that p/q > 1/2?

jhonyy9 (jhonyy9):

11/20

OpenStudy (jamesj):

I ask because there is no smallest fraction 1/2 < 51/100 < 11/20 But then 1/2 < 501/1000 < 51/100 But then 1/2 < 5001/10000 < 501/1000 < 51/100 < 11/20

OpenStudy (jamesj):

Given any fraction greater than 1/2, you can always find another fraction between them.

OpenStudy (jamesj):

..for example, the fraction half way between them: \[ \frac{1}{2} \left( \frac{p}{q} + \frac{1}{2} \right) =\frac{1}{2} \frac{ 2p + q}{ 2q } = \frac{ 2p + q}{ 4q } \] So suppose you started with 11/20. The fraction half way between 1/2 and 11/20 (p = 11, q = 20) is \[ \frac{1}{2} \left( \frac{11}{20} + \frac{1}{2} \right) = \frac{22 + 20}{80} = \frac{21}{40} \]